HONG KONG DIPLOMA OF SECONDARY EDUCATION EXAMINATION 2027 · PRACTICE PAPER
MATHEMATICS Extended PartModule 1 (Calculus and Statistics)
Question-Answer Book
8.30 am – 11.00 am (2½ hours) This paper must be answered in English
INSTRUCTIONS
(1)After the announcement of the start of the examination, you should first write your Candidate Number in the space provided on Page 1 and stick barcode labels in the spaces provided.
(2)This paper consists of TWO sections, A and B.
(3)Attempt ALL questions in this paper. Write your answers in the spaces provided in this Question-Answer Book. Do not write in the margins. Answers written in the margins will not be marked.
(4)Graph paper and supplementary answer sheets will be supplied on request. Write your Candidate Number, mark the question number box and stick a barcode label on each sheet, and fasten them with string INSIDE this book.
(5)Unless otherwise specified, all working must be clearly shown.
(6)Unless otherwise specified, numerical answers should be either exact or given to 4 decimal places.
(7)No extra time will be given to candidates for sticking on the barcode labels or filling in the question number boxes after the 'Time is up' announcement.
Studify practice paper. Not an official HKEAA examination paper.
Please stick the barcode label here.
Candidate Number
Answers written in the margins will not be marked.
Answers written in the margins will not be marked.
SECTION A (50 marks)
1.
The table below shows the probability distribution of a discrete random variable X, where a and b are constants. It is given that E(X)=3.
(a) Find a and b.
(b) Find Var(X) and Var(3−2X).(6 marks)
x
1
2
4
7
P(X = x)
0.2
a
b
0.1
Answers written in the margins will not be marked.
Answers written in the margins will not be marked.
2.
Let A and B be two events. Denote the complementary events of A and B by A′ and B′ respectively. It is given that P(A)=0.4, P(B∣A)=0.3 and P(A∪B)=0.7.
(a) Find P(B).
(b) Are A and B independent? Explain your answer.
(c) Find P(A′∣B′).(5 marks)
Answers written in the margins will not be marked.
Answers written in the margins will not be marked.
3.
The number of customers arriving at a convenience store in a 10-minute interval follows a Poisson distribution with a mean of 1.8.
(a) Find the probability that at least 2 customers arrive at the store in a 10-minute interval.
(b) Find the probability that exactly 3 customers arrive at the store in a 20-minute interval.
(c) Find the probability that, in 5 separate 10-minute intervals, there are exactly 3 intervals in each of which at least 2 customers arrive at the store.(7 marks)
Answers written in the margins will not be marked.
Answers written in the margins will not be marked.
4.
The weights of apples from an orchard are normally distributed with a standard deviation of 15 g. A random sample of 36 apples is taken and their mean weight is 182 g.
(a) Construct a 95% confidence interval for the mean weight of the apples from the orchard.
(b) Find the least sample size such that the width of a 99% confidence interval for the mean weight is less than 8 g.(6 marks)
Answers written in the margins will not be marked.
Answers written in the margins will not be marked.
5.
(a) Expand e3x in ascending powers of x as far as the term in x3.
(b) Find the coefficient of x3 in the expansion of e3x(1−2x)5.(6 marks)
Answers written in the margins will not be marked.
Answers written in the margins will not be marked.
6.
Consider the curve C:y=ln(2x−3)+x2, where x>23.
(a) Find dxdy.
(b) Find the equation of the tangent to C at the point where x=2.
(c) The tangent in (b) cuts the x-axis and the y-axis at P and Q respectively. Let O be the origin. Find the area of △OPQ.(7 marks)
Answers written in the margins will not be marked.
Answers written in the margins will not be marked.
7.
Let g(x) be a function such that g′(x)=x2+14x for all real numbers x. It is given that g(0)=3.
(a) Find g(x).
(b) Find the exact value of the area of the region bounded by the curve y=g′(x), the x-axis and the straight line x=2.(6 marks)
Answers written in the margins will not be marked.
Answers written in the margins will not be marked.
8.
The profit P (in million dollars) of a company t years after its establishment is modelled by P=20te−0.25t, where 0≤t≤10.
(a) Find dtdP.
(b) Find the greatest profit of the company and the value of t at which it occurs.
(c) Is the profit increasing when t=6? Explain your answer.(7 marks)
Answers written in the margins will not be marked.
Answers written in the margins will not be marked.
SECTION B (50 marks)
9.
A factory has two machines, A and B, which produce 60% and 40% of its chips respectively. 2% of the chips produced by machine A and 5% of the chips produced by machine B are defective.
(a)
Find the probability that a randomly selected chip produced by the factory is defective.(1 分)
(b)
Given that a randomly selected chip is defective, find the probability that it is produced by machine B.(2 分)
(c)
The chips are packed in boxes of 20. A box is accepted if it contains at most 1 defective chip.
(i) Find the probability that a box is accepted.
(ii) An inspector checks boxes one by one. Find the probability that the first box which is not accepted is the 4th box checked.
(iii) Find the probability that exactly 8 out of 10 boxes are accepted.(9 分)
Answers written in the margins will not be marked.
Answers written in the margins will not be marked.
Answers written in the margins will not be marked.
Answers written in the margins will not be marked.
10.
The daily commuting time of the workers in a city follows a normal distribution with a mean of 42 minutes and a standard deviation of 8 minutes.
Extract from the table of the area under the standard normal curve between 0 and z
z
1.00
1.28
1.50
1.96
Area
.3413
.3997
.4332
.4750
(a)
Find the probability that the daily commuting time of a randomly selected worker exceeds 50 minutes.(2 分)
(b)
Find t such that 10% of the workers have daily commuting times of more than t minutes.(2 分)
(c)
16 workers are randomly selected.
(i) Find the probability that the mean daily commuting time of the 16 workers exceeds 45 minutes.
(ii) Find the probability that exactly 3 of the 16 workers have daily commuting times of more than 50 minutes.(4 分)
(d)
A year later, a random sample of 64 workers has a mean daily commuting time of 44.5 minutes. Assume that the standard deviation is still 8 minutes.
(i) Construct a 95% confidence interval for the mean daily commuting time μ minutes of the workers.
(ii) A reporter claims that the mean daily commuting time of the workers has increased. Using the result of (d)(i), do you agree? Explain your answer.(5 分)
Answers written in the margins will not be marked.
Answers written in the margins will not be marked.
Answers written in the margins will not be marked.
Answers written in the margins will not be marked.
Answers written in the margins will not be marked.
Answers written in the margins will not be marked.
11.
Let f(x)=ln(x2+1). Denote the region bounded by the curve y=f(x), the x-axis and the straight lines x=1 and x=3 by R.
(a)
Using the trapezoidal rule with 4 sub-intervals, estimate ∫13f(x)dx.(3 分)
(b)
Is the estimate in (a) an over-estimate or an under-estimate? Explain your answer.(3 分)
(c)
Someone claims that the area of R exceeds 3.1. Do you agree? Explain your answer.(2 分)
(d)
(i) Find the equation of the tangent to the curve y=f(x) at the point where x=2.
(ii) Hence, show that the area of R is less than 2ln5.(4 分)
Answers written in the margins will not be marked.
Answers written in the margins will not be marked.
Answers written in the margins will not be marked.
Answers written in the margins will not be marked.
12.
A fisheries officer models the number of fish N (in thousands) in a reservoir by N=1+4e−0.5t50, where t is the number of years elapsed since the start of the study.
(a)
Find the number of fish when t=0.(1 分)
(b)
Express ln(N50−1) as a linear function of t, and write down the slope and the intercept on the vertical axis of its graph.(2 分)
(c)
Find dtdN.(2 分)
(d)
Find the value of t at which the rate of increase of the number of fish is the greatest.(4 分)
(e)
Estimate the number of fish in the reservoir after a very long time.(1 分)
(f)
The officer claims that the number of fish will exceed 45 000 within the first 6 years of the study. Do you agree? Explain your answer.(3 分)
Answers written in the margins will not be marked.
Answers written in the margins will not be marked.
(a) P(A∩B)=0.4(0.3)=0.12 1M
P(B)=0.7−0.4+0.12=0.42 1A
(b) P(A)P(B)=0.168=0.12=P(A∩B), so A and B are not independent. 1A f.t.
(c) P(A′∣B′)=1−P(B)1−P(A∪B)=0.580.3 1M
=2915≈0.5172 1A (5)
第 3 題7 分
(a) 1−e−1.8(1+1.8) 1M
≈0.5372 1A
(b) The number in a 20-minute interval follows a Poisson distribution with mean 3.6. 1M
3!e−3.6(3.6)3≈0.2125 1A
(c) C35(0.537163)3(1−0.537163)2 1M+1M
≈0.3320 1A (7)
第 4 題6 分
(a) 182±1.96(3615) 1M
The confidence interval is (177.1,186.9). 1A
(b) 2(2.576)(n15)<8 1M+1M
n>93.3 1A
The least sample size is 94. 1A (6)
(Accept 2.575 for the critical value.)
(a) dxdy=2x−32+2x 1A
(b) When x=2, y=ln1+4=4 and dxdy=2+4=6. 1M
The equation of the tangent is y−4=6(x−2), i.e. y=6x−8. 1A
(c) P=(34,0) and Q=(0,−8) 1M+1A
Area of △OPQ=21(34)(8) 1M
=316 1A (7)
第 7 題6 分
(a) Let u=x2+1. g(x)=∫x2+14xdx=∫u2du 1M
=2ln(x2+1)+c 1A
g(0)=3 gives c=3, so g(x)=2ln(x2+1)+3. 1A
(b) g′(x)≥0 for x≥0, so the area is ∫02x2+14xdx 1M
=[2ln(x2+1)]02 1M
=2ln5 1A (6)
第 8 題7 分
(a) dtdP=20e−0.25t+20t(−0.25)e−0.25t=20e−0.25t(1−0.25t) 1M+1A
(b) dtdP=0 gives t=4. 1M
dtdP>0 for 0≤t<4 and dtdP<0 for 4<t≤10, so P is greatest when t=4. 1M
Greatest profit =80e−1≈29.4304 million dollars 1A
(c) When t=6, dtdP=20e−1.5(1−1.5)<0 1M
So the profit is not increasing. 1A f.t. (7)
(a) P(X>50)=P(Z>1)=0.5−0.3413=0.1587 1M+1A
(b) P(Z>z)=0.1 gives z≈1.28, so t≈42+1.28(8)=52.24 1M+1A
(c)
(i) Xˉ∼N(42,1682), so P(Xˉ>45)=P(Z>1.5) 1M
=0.0668 1A
(ii) C316(0.1587)3(0.8413)13 1M
≈0.2367 1A
(d)
(i) 44.5±1.96(648) 1M+1M
The confidence interval is (42.54,46.46). 1A
(ii) 42 is not in the confidence interval in (d)(i), 1M
so there is evidence that the mean commuting time has increased. The claim is agreed. 1A f.t. (13)
第 11 題12 分
(a) ∫13f(x)dx≈20.5[ln2+2(ln3.25+ln5+ln7.25)+ln10] 1M+1M
≈3.1335 1A
(b) f′(x)=x2+12x 1A
f′′(x)=(x2+1)22(x2+1)−2x(2x)=(x2+1)22(1−x2) 1M
For 1<x<3, f′′(x)<0, so the curve is concave downward and the estimate in
(a) is an under-estimate. 1A f.t.
(c) Area of R>3.1335>3.1, so the claim is agreed. 1M+1A f.t.
(d)
(i) f(2)=ln5 and f′(2)=54, so the tangent is y=ln5+54(x−2). 1A
(ii) Since the curve is concave downward on 1≤x≤3, it lies below the tangent there. 1M
Area of R<∫13[ln5+54(x−2)]dx=2ln5+54[2(x−2)2]13=2ln5 1M+1A (12)
第 12 題13 分
(a) N=1+450=10 1A
(b) N50−1=4e−0.5t, so ln(N50−1)=ln4−0.5t 1M
The slope is −0.5 and the intercept on the vertical axis is ln4. 1A
(c) dtdN=(1+4e−0.5t)2−50(4)(−0.5)e−0.5t=(1+4e−0.5t)2100e−0.5t 1M+1A
(d) Let u=e−0.5t. dtdN=(1+4u)2100u and dud[(1+4u)2100u]=(1+4u)3100(1−4u) 1M
This is 0 when u=41; it is positive for u<41 and negative for u>41. 1M
e−0.5t=41 gives t=2ln4≈2.7726 1A
Since u decreases as t increases, dtdN is greatest at t=2ln4. 1A
(e) As t→∞, e−0.5t→0, so N→50, i.e. about 50 000 fish. 1A
(f) When t=6, N=1+4e−350≈41.6963 1M
N is increasing for all t, so N<41.6963<45 for 0≤t≤6. 1M
The claim is disagreed. 1A f.t. (13)